# Depth Matters: Multimodal RGB-D Perception for Robust Autonomous Agents

### v1

@@ Abstract
Autonomous agents that rely purely on perception to make real-time control decisions require efficient and robust architectures. In this work, we demonstrate that augmenting RGB input with depth information significantly enhances our agents’ ability to predict steering commands compared to using RGB alone. We benchmark lightweight recurrent controllers that leverage the fused RGB-D features for sequential decision-making. To train our models, we collect high-quality data using a small-scale autonomous car controlled by an expert driver via a physical steering wheel, capturing varying levels of steering difficulty. Our models, trained under diverse configurations, were successfully deployed on real hardware. Specifically, our findings reveal that the early fusion of depth data results in a highly robust controller, which remains effective even with frame drops and increased noise levels, without compromising the network’s focus on the task.
@@ I INTRODUCTION
On the one hand, unimodal RGB input has been successfully used in real-time applications in prior work [1], where a hybrid architecture of convolutional-head with biologically inspired Liquid Time-Constant (LTC) [2] networks was used. They demonstrated exceptional performance in autonomous lane-keeping tasks by directly mapping raw RGB camera inputs to steering commands, outperforming conventional deep networks in both interpretability and resilience to sensor noise. One of the goals is to replicate the model architecture used here as closely as possible, with the necessary adjustments. However, it is known that differential equation (DE) solvers used in their model can cause longer computation times. In recent work, [3] introduced Closed-form Continuous-time Neural Network (CfC), a variant of the ODE-based continuous LTC network. Another efficient extension is by introducing a liquid capacitance term into the model [4] on top of saturation [5], called Liquid Resistance Liquid Capacitance Network (LRC). These solutions can be beneficial in more lightweight hardware setups, ensuring faster reaction times, which we implement and test in this paper.
Figure 1: High-level overview of the model architectures deployed on our hardware, following a channels-first configuration
On the other hand, unimodal approaches face inherent limitations in complex real-world scenarios where depth perception is critical for obstacle avoidance and path planning. To tackle this challenge, we employ multimodal sensory fusion. RGB-D input is used in simulation both in [6] and  [7]. Although these approaches demonstrate success in simulated environments, their applicability to real-world scenarios remains unexplored, especially on small-scale platforms like roboracer [8]. This paper seeks to bridge that gap through its real-world application.
In addition, the fusion of RGB-D information has been widely researched in computer vision, mainly using Convolutional Neural Networks (CNN)[9]. A widely used approach is introducing depth information as a new channel to the input of 2D convolutions, which yields the best results in [7]. Alternatively, the authors of [10] use depth information in the computation of Deformable CNNs [11, 12], which aim to enhance spatial flexibility. Their approach leads to more accurate object classification than with RGB alone. We aim to test these more advanced approaches not in computer vision but in a control task.
This paper aims to investigate the integration of RGB-D data for autonomous driving in a lightweight real-world scenario. In this way, we can achieve optimized multimodal models for the small-scale autonomous system roboracer with fast reaction times and behavior comparable to that of an experienced human driver.
@@
  • We show that depth information is crucial, RGB alone is not enough for robust navigation.
  • A systematic comparison of RGB-D fusion architectures for autonomous driving, evaluating depth-as-channel (early fusion and late fusion) and depth-aware deformable CNNs (DCN [11] and ZACN[13]) to optimize feature extraction for robust perception in noisy real-world environments.
  • Lightweight recurrent controller benchmarking for real-time efficiency, which uses the extracted features from our multimodal vision block from the previous point for control.
  • Bridging the sim-to-real gap for RGB-D multimodal models with recurrent controllers: the first deployment of a depth-enhanced network on a physical roboracer vehicle, achieving a smooth and noise-robust controller.
The rest of the paper is organized as follows. Section II reviews prior work on multimodal perception and autonomous driving architectures. In Section III, we describe our hardware configuration and dataset collection process. Section IV summarizes our proposed model architectures integrating RGB-D fusion. In Section V, we evaluate the impact of depth-aware feature extraction via vision fusion heads paired with recurrent controllers for sequential steering prediction. Finally, we discuss our results in Section VI.
@@ II-A Unimodal RGB Input
A critical achievement in the context of end-to-end autonomous driving is [14], in which a 9-layer Convolutional Neural Network (CNN) trained on RGB input successfully predicted steering commands on a drive-by-wire vehicle. Although it is shown the model learned useful road features, it remained impossible to assess which segments of the network contributed to its decisions due to the large scale of the network.
Alternatively, [15] tackled the limitations of traditional CNN-based steering models by introducing a temporal dependency-aware approach. They propose a C-LSTM network architecture with CNN for feature extraction, followed by Long Short-Term Memory (LSTM) [16]. After training with RGB images obtained from a simulated environment, the evaluation of the model revealed effective feature extraction and autonomy in mostly obstacle-free scenarios. However, real-world deployment remained untested.
Addressing the issues of interpretability and real-time use that such traditional deep learning methods pose, the authors of [1] introduced a compact bio-inspired neural controller, called Neural Circuit Policy (NCP). Their model significantly out-performed purely CNN methods both in simulation and hardware deployment, using RGB data.
Another study incorporates optical flow extracted from RGB frames to improve steering estimation [17]. Their approach integrates CNN-based feature extraction with Recurrent Neural Network (RNN)’s, including LSTM and NCP. By employing both early and hybrid fusion techniques, they demonstrated that incorporating optical flow significantly reduces steering estimation error compared to RGB-only methods in dataset-only evaluation. Nevertheless, the results are limited to dataset experiments, without on-board evaluation.
@@ II-B Multimodal Input
Even though several studies have explored different approaches to improve the performance of autonomous driving models by leveraging RGB-D information, many of them lack real-world testing to confirm practical viability.
Notably, [6] presents a Recurrent CNN (RCNN) trained by fusing camera and LiDAR data. Their results accentuate the importance of multimodal perception for enhancing autonomy, outperforming RGB-only models. However, despite its strong performance in simulation, the model was not evaluated in a real-world driving scenario, leaving its practical deployment untested.
Similarly, another study investigates the impact of multimodal perception on fully autonomous driving, particularly focusing on early, mid, and late fusion techniques [7]. Using the CARLA simulator, the authors compare the performance of single-modality (RGB/depth) and multimodality. Their results indicate that early fusion multimodality significantly outperforms single-modality models, especially under challenging driving conditions. However, real-world validation was not performed, limiting its practical insights.
II-C State-of-the-Art of roboracer
End-to-end (E2E) deep learning approaches have been widely explored for autonomous racing on the roboracer platform, with multiple works leveraging Deep Reinforcement Learning (DRL) for control [18, 19, 20]. In contrast, supervised learning-based approaches for roboracer are relatively unexplored, with TinyLidarNet [21] being a key recent contribution. TinyLidarNet introduces a lightweight LiDAR-based CNN model, which maintains fast inference efficiency on low-end microcontrollers. However, its reliance on LiDAR raises concerns about robustness in challenging environments. As a separate study [22] has shown, LiDAR performance drops to 33% while depth cameras maintain full-ranging capabilities in adverse conditions. Given these findings, our work diverges from TinyLidarNet by adopting an RGB-D-based supervised learning approach, aiming to benefit from stereoscopic camera sensing’s robustness and resilience to lighting variations.
@@ III METHODS
In this section, we present the methodology used to develop and evaluate our model architectures. We start by describing the hardware setup and our data collection process, followed by outlining the dataset creation pipeline, including the measures implemented to ensure high-quality training data.
@@ III-A Hardware Setup and Data Collection
The platform we used in this paper is a one-tenth-scaled car, referred to as roboracer in the autonomous driving research community. Further details about the roboracer platform, including in-depth hardware description and state-of-the-art applications are detailed by the platform creators in [23].
The major hardware components of the roboracer platform include a lower-level chassis, an upper-level chassis, and several autonomy elements. The embedded computing platform is a 13th Gen Intel(R) Core(TM) i7-1360P. Most importantly for optical vision, we integrated a stereoscopic depth camera (Intel RealSense D435i) into the hardware stack. According to the datasheet, key application-specific properties of the camera include a range of 10m (best performance between 0.3m and 3m), and 69°$\degree$$\times$42°$\degree$ RGB FOV, 87°$\degree$$\times$58°$\degree$ Depth FOV.
Given the 30fps limitation of the RGB camera, the same number of fps was used for the depth stream. Furthermore, both camera streams were configured to 848$\times$480px resolution, which is the optimal recommended resolution for the depth camera.
@@
We built a series of 5 tracks to offer a variation of successive turns, either of different curvatures or lengths before, during, and after each turn, as seen in Figure 2. In order to increase the dataset size, all tracks were driven both clockwise and anti-clockwise.
A driver with 10 years of frequent driving experience was chosen to control the car on different tracks. The driver was given control of the vehicle through a racing steering wheel and pedals, with a speed cap of 0.9 m/s. We introduced the speed cap after several trials were made by the human driver around track Figure 2(a) without any speed restrictions. We found that due to the narrow FOV of the camera, the driver was most comfortable driving with a maximum speed of 0.9m/s.
The data flow of the driving stack during data collection is illustrated in Figure 3. In order to simulate a first-person point of view, we redirected the RGB stream from the onboard camera via the network to a second device with a monitor placed in front of the driver. While driving, rosbags were being recorded on all available car topics, excluding the RGB topic. The RGB data stream was saved as videos while driving on the support device connected to the monitor.
@@
Figure 3: Overview of our driving stack used during the data collection step.
The driver was allowed a few trial runs on each track before the data recording was started in order to get accustomed to the camera’s field of view and the difficulty of the track. For Figure 2(a), we instructed the driver to take the longer route at the intersection.
Figure 4: The sequence of layers in the convolutional heads is as follows. The Convolutional Feature Extractor is used with an RGB input of size 3×120×212. In the early fusion (EARLY) method, it processes an input of size 4×120×212. In the late fusion (LATE) method, it is used twice, treating RGB and depth as two separate input streams of sizes 3×120×212 and 1×120×212, respectively, and concatenating their features after the last layer. The Depth-Adapted Convolutional Feature Extractor has two versions: (1) the DCN, which includes a Convolutional Offset Extractor, and (2) the ZACN [13], which incorporates Geometric Offset Computation. In the hyperparameters section: F denotes the number of filters, K the kernel size, S the stride, and P the padding.
@@ III-B Dataset Collected for Open-Loop Training
To synchronize the rosbags and RGB frames, we leveraged the mutual 30fps rate chosen to configure the camera.
Firstly, the messages on the depth topic were extracted from each rosbag [24]. Their content consists of uint16 1-channel images representing the measured distance for each pixel. We parsed each frame to RGB by creating a colormap on the pixels based on their distance with respect to the maximum measured value. Then, we generated depth videos from each rosbag and stamped every frame with its corresponding number.
As a second step, we manually synchronized the original videos with the depth. The original videos were also broken down and stamped by frame number as a preparatory procedure. Before the start of each recording session, we presented visual cues to the camera to synchronize the streams. In this way, we were able to log the frame number corresponding to each visual cue for both the RGB and depth streams. Furthermore, we logged frame numbers corresponding to the start and end of every lap in all recordings.
Thirdly, the resolution of the streams was downscaled to 1/4th of the initial scale. For this purpose, we tested several rescaling algorithms: nearest neighbor, Gaussian, Lanczos [25], and Catmull-Rom [26]. As expected, nearest neighbor downsampling produced grainier results, while Gaussian downsampling blurred the stream. As neither was desirable, Lanczos-3 was chosen over Catmull-Rom for producing overall sharper results.
Lastly, the depth topic messages were iterated over based on the lap frame intervals in each recording. This led to obtaining the timestamp of each frame message, which we then used to find the corresponding steering commands as in [24] by interpolating the values between two timestamps. In this way, each RGB frame and depth message was labeled with the commanded steering angle and saved as .npz files. Through this approach, we obtained a dataset of approximately 100 000100000100\,000100 000 labeled frames from all collected data.
@@ III-C Work Environment
The machine used for training the models runs Ubuntu 20.04.6 and has 3 07230723\,0723 072 CUDA cores and 12 GB of GDDR5 memory. Due to its out-of-the-box interoperability and wide range of packages in the area of computer vision, we chose PyTorch as a framework, and configured a Conda virtual environment. Code is available in our GitHub repository.
@@ IV-A Unimodal RGB Setup
The implemented unimodal architecture closely follows [1], with a convolutional head illustrated in Figure 4 consisting of 6 convolutional layers which feed their final output to the RNN of choice. This approach aims to extract relevant features of the RGB input and use them as input to RNN s, which make memory-based steering commands. We refer to this method as RGB in the next sections.
@@ IV-B Multimodal RGB-D Setups
(i) The first multimodal approach is an early fusion approach (referred to as EARLY), introducing the depth values as a fourth channel to the RGB input of the previously described convolutional head. This type of method for CNN’s achieved the best results in [6] and [7]. This architecture is shown in Figure 4 as the Convolutional Feature Extractor.
(ii) Similarly to [27], the second multimodal architectural method is a late fusion method (referred to as LATE). This essentially presents a two-stream architecture, where different CNN’s are used for each of the inputs.
(iii) The final chosen approach illustrated in Figure 4 as Depth-Adapted Convolutional Feature Extractor, is processing the multimodal input using Deformable CNN. As the goal of deformable convolutions is to hone spatial adaptation of kernels, we only use the operation once, in the last layer. In this way, we ensure that the most relevant features extracted from RGB using standard convolution are simply enhanced by offsets that are extracted from depth information. To this end, two techniques were utilized for calculating the convolutional offsets for the deformable layer.
@@
(iii-b) The second technique, which we refer to as ZACN, is based on [13]. The intrinsic camera parameters used by this method were obtained from the device used during data collection.

### v3

@@ Abstract
Autonomous agents that rely purely on perception to make real-time control decisions require efficient and robust architectures. In this work, we demonstrate that augmenting RGB input with depth information significantly enhances our agents’ ability to predict steering commands compared to using RGB alone. We benchmark lightweight recurrent controllers that leverage the fused RGB-D features for sequential decision-making. To train our models, we collect high-quality data using a small-scale autonomous car controlled by an expert driver via a physical steering wheel, capturing varying levels of steering difficulty. Our models were successfully deployed on real hardware and inherently avoided dynamic and static obstacles, under out-of-distribution conditions. Specifically, our findings reveal that the early fusion of depth data results in a highly robust controller, which remains effective even with frame drops and increased noise levels, without compromising the network’s focus on the task.
@@ I INTRODUCTION
On the one hand, unimodal RGB input has been successfully used in real-time applications in prior work [Lechner_Hasani_Amini_Henzinger_Rus_Grosu_2020], where a hybrid architecture of convolutional-head with biologically inspired Liquid Time-Constant (LTC) [hasani2021liquid] networks was used. They demonstrated exceptional performance in autonomous lane-keeping tasks by directly mapping raw RGB camera inputs to steering commands, outperforming conventional deep networks in both interpretability and resilience to sensor noise. One of the goals is to replicate the model architecture used here as closely as possible, with the necessary adjustments. However, it is known that differential equation (DE) solvers used in their model can cause longer computation times. In recent work, [Hasani_2022] introduced Closed-form Continuous-time Neural Network (CfC), a variant of the ODE-based continuous LTC network. Another efficient extension is by introducing a liquid capacitance term into the model [farsang2024liquid] on top of saturation [farsang2024learning], called Liquid Resistance Liquid Capacitance Network (LRC). These solutions can be beneficial in lightweight hardware setups, ensuring faster reaction times, which we implement and test in this paper.
Figure 1: High-level overview of the model architectures.
On the other hand, unimodal approaches face inherent limitations in complex scenarios where depth perception is critical for obstacle avoidance and path planning. To tackle this challenge, we employ multimodal sensory fusion. RGB-D input is used in simulation both in [Mújica-Vargas2024] and  [MultimodalEnd-to-EndAutonomousDriving]. Although such approaches demonstrate success in simulated environments, their applicability to real-world scenarios remains unexplored, especially on small-scale platforms like roboracer [f1tenth]. This paper seeks to bridge that gap, both as a real-world application, and with more lightweight architectures.
In addition, the fusion of RGB-D information has been widely researched in computer vision, mainly using Convolutional Neural Networks (CNN)[lecuncnn]. A widely used approach is introducing depth information as a new channel to the input of 2D convolutions, which yields the best results in [MultimodalEnd-to-EndAutonomousDriving]. Alternatively, the authors of [depthawarecnn] use depth information in the computation of Deformable CNNs [deformcnn, zhu2018deformableconvnetsv2deformable], which aim to enhance spatial flexibility. Their approach leads to more accurate object classification than with RGB alone. We aim to test these more advanced approaches not in computer vision but in a navigation task.
This paper aims to investigate the use of RGB-D data for navigating in complex scenarios. In this way, we can achieve optimized multimodal models for the small-scale autonomous system roboracer with fast reaction times and behavior comparable to that of an experienced human driver.
@@
  • We show depth information is crucial, RGB alone is not enough for robust navigation, as RGB-only agents fail in scenarios with obstacles, decision points, or sharp turns while RGB-D agents succeed.
  • A systematic comparison of RGB-D fusion techniques, evaluating depth-as-channel (early, late fusion) and depth-aware deformable CNNs (DCN [deformcnn] and ZACN[wu2020depthadaptedcnnrgbdcameras]) to optimize feature extraction for perception.
  • Lightweight recurrent controller benchmarking for efficiency, using features extracted from our multimodal vision block from the previous point for control.
  • The first deployment of a depth-enhanced network on a roboracer vehicle, achieving a smooth and noise-robust control and zero-shot adaptation to unseen scenarios such as static and dynamic obstacles and intersections.
@@ II-A Unimodal RGB Input
A critical achievement in the context of end-to-end autonomous driving is [DBLP:journals/corr/BojarskiTDFFGJM16], in which a Convolutional Neural Network (CNN) trained on RGB input successfully predicted steering commands on a drive-by-wire vehicle. Although it is shown the model learned useful road features, it remained impossible to assess which segments of the network contributed to its decisions due to the large scale of the network.
Alternatively, [9144508] tackled the limitations of traditional CNN-based steering models by adding temporal awareness via Long Short-Term Memory (LSTM) [lstm]. After training on synthetic RGB images, the model extracted features effectively and navigated obstacle-free scenarios, but real-world deployment was not tested.
Addressing the issues of interpretability of traditional deep learning methods, the authors of [Lechner_Hasani_Amini_Henzinger_Rus_Grosu_2020] introduced a compact bio-inspired neural controller, called Neural Circuit Policy (NCP). Their model significantly outperformed purely CNN methods in hardware deployment, using RGB data.
Another study extracts optical flow from RGB frames to improve steering estimation [makiyeh2024opticalflowmattersempirical]. By employing both early and hybrid fusion techniques with CNN and Recurrent Neural Network (RNN)’s, they demonstrated that incorporating optical flow significantly reduces steering estimation error compared to RGB-only methods in dataset-only evaluation. Nevertheless, the results are limited to dataset experiments, without on-board evaluation.
@@ II-B Multimodal Input
Although several studies have explored different approaches to improve the performance of autonomous driving models by leveraging RGB-D information, many of them lack real-world testing to confirm practical viability.
Notably, [Mújica-Vargas2024] presents a large CNN-LSTM trained by fusing camera and LiDAR data. Their results accentuate the importance of multimodal perception for enhancing autonomy, outperforming RGB-only models. However, the model was not evaluated in a real-world driving scenario, leaving its practical deployment untested.
Similarly, another study investigates the impact of multimodal perception wit a branched CNN-MLP architecture, focusing on fusion techniques [MultimodalEnd-to-EndAutonomousDriving]. Using the CARLA simulator, the authors compare the performance of single-modality (RGB/depth) and multimodality. Their results indicate that early fusion multimodality outperforms single-modality models. However, real-world validation was not performed, limiting its practical insights.
II-C roboracer: Learning Approaches and Sensor Modalities
End-to-end (E2E) deep learning approaches have been widely explored using the roboracer platform, with multiple works leveraging Deep Reinforcement Learning (DRL) for control [evans2023bypassingsimulationtorealitygaponline, EVANS2023107, EVANS2023100496]. However, most of these studies rely on the available simulators: the F1Tenth 2D simulator [pmlr-v123-o-kelly20a] or the ROS/Gazebo-based variant [9216949], which do not provide realistic rendering for RGB-D cameras. In particular, the Gazebo-based camera models fail to capture the visual complexity of real-world environments, making it difficult to evaluate complex pipelines in simulation. This motivates the need for real-world data when studying robotic controllers. In contrast, supervised learning-based approaches for roboracer are relatively unexplored, with TinyLidarNet [zarrar2024tinylidarnet2dlidarbasedendtoend] being a recent contribution. TinyLidarNet introduces a lightweight LiDAR-based CNN model, which maintains fast inference efficiency on low-end microcontrollers. However, its reliance on LiDAR raises concerns about robustness in challenging environments. As a separate study [loetscher2023assessingrobustnesslidarradar] has shown, LiDAR performance drops to 33% while depth cameras maintain full-ranging capabilities in adverse conditions. Given these findings, our work diverges from TinyLidarNet by adopting an RGB-D-based supervised learning approach, aiming to benefit from stereoscopic camera sensing’s robustness and resilience to lighting variations. Moreover, to provide a fair comparison and to validate our sensor choice, we also trained and deployed TinyLidarNet under the same experimental conditions as our RGB-D models. This allowed us to directly contrast the two sensing modalities and demonstrate the robustness of RGB-D in scenarios with reflective surfaces.
@@ III METHODS
In this section, we present the methodology used to develop and evaluate our neural architectures. We start by describing the hardware setup and our data collection process, followed by the dataset creation pipeline, including the measures implemented to ensure high-quality training data.
@@ III-A Hardware Setup and Data Collection
The platform we used in this paper is a 1/10-scaled car, referred to as roboracer in the autonomous driving research community. Further details about the platform, including in-depth hardware description and state-of-the-art applications are detailed by the platform creators in [f1tenthcourse].
[…]
Given the 30fps limitation of the RGB camera, the same number of fps was used for the depth stream. Both camera streams were configured to $8×\times4$80px resolution, which is the optimal recommended resolution for the depth camera.
@@
We built a series of 5 tracks with a variation of successive turns, of different curvatures and lengths, as seen in Figure 2. In order to increase the dataset size, all tracks were driven both clockwise and anti-clockwise.
A driver with 10 years of driving experience was chosen to control the car. The driver was given control of the vehicle through a racing steering wheel and pedals. We found that due to the narrow FOV of the camera, the driver was most comfortable driving with a maximum speed of 0.9m/s.
The data flow of the driving stack during data collection is illustrated in Figure 3. To simulate a first-person point of view, we redirected the RGB stream from the camera to a second device with a monitor placed in front of the driver. While driving, rosbags were being recorded on all available car topics, excluding the RGB topic, since the RGB data stream was being saved as video files.
@@
Figure 3: Driving stack overview for data collection step.
Figure 4: The sequence of layers in the convolutional heads is as follows. The Convolutional Feature Extractor is used with an RGB input of size 3×120×212. In the early fusion (EARLY) method, it processes an input of size 4×120×212. In the late fusion (LATE) method, it is used twice, treating RGB and depth as two separate input streams of sizes 3×120×212 and 1×120×212, respectively, and concatenating their features after the last layer. The Depth-Adapted Convolutional Feature Extractor has two versions: (1) the DCN, which includes a Convolutional Offset Extractor, and (2) the ZACN [wu2020depthadaptedcnnrgbdcameras], which incorporates Geometric Offset Computation. In the hyperparameters section: F denotes the number of filters, K the kernel size, S the stride, and P the padding.
@@ III-B Dataset Collected for Open-Loop Training
Firstly, the messages on the depth topic were extracted from each rosbag. Their content consists of uint16 1-channel images representing the measured distance for each pixel. We parsed each frame to RGB by creating a colormap on the pixels based on their distance with respect to the maximum measured value. Then, we generated depth videos from each rosbag and stamped every frame with its index.
As a second step, we manually synchronized the original videos with the depth from the rosbags. The original videos were broken down and stamped by frame number. Before the start of each recording session, we presented visual cues to the camera to help with synchronization. Furthermore, we logged frame numbers corresponding to the start and end of each lap in the recordings.
Thirdly, the resolution of the streams was downscaled to 1/4th of the initial scale. For this purpose, we tested several rescaling algorithms: nearest neighbor, Gaussian, Lanczos [lanczos1950iteration], and Catmull-Rom [catmull1974class]. As expected, nearest neighbor downsampling produced grainier results, while Gaussian downsampling blurred the stream. As neither was desirable, Lanczos-3 was chosen over Catmull-Rom for producing overall sharper results.
Lastly, the depth topic messages were iterated over based on the lap frame intervals in each recording. With the timestamp of each frame message, we then found the corresponding steering commands as in [resch2023attention] by interpolating the values between two timestamps. In this way, each RGB frame and depth message was labeled with the commanded steering angle and saved as .npz files. We obtained a dataset of approximately 100 000100\,000 labeled frames.
@@ III-C Work Environment
The machine used for training the models runs Ubuntu 20.04.6 and has 3 0723\,072 CUDA cores and 12 GB of GDDR5 memory. We chose PyTorch due to its out-of-the-box interoperability and extensive computer vision package support.
@@ IV-A Unimodal RGB Setup
The implemented unimodal architecture follows [Lechner_Hasani_Amini_Henzinger_Rus_Grosu_2020], with a convolutional head illustrated in Figure 4 which feeds the output to an RNN of choice. This aims to extract relevant features of the RGB input and use them as input to RNN s to make memory-based steering commands. We refer to this method as RGB in the next sections.
@@ IV-B Multimodal RGB-D Setups
(i) The first multimodal approach is an early fusion approach (referred to as EARLY), introducing the depth values as a fourth channel to the RGB input of the previously described convolutional head. This type of fusion achieved the best results in [Mújica-Vargas2024] and [MultimodalEnd-to-EndAutonomousDriving]. Our specific architecture is shown in Figure 4 as the Convolutional Feature Extractor.
(ii) The second multimodal method is a late fusion method (referred to as LATE). This essentially presents a two-stream architecture, where different CNN’s are used for each stream.
(iii) The final method illustrated in Figure 4 as Depth-Adapted Convolutional Feature Extractor, is processing the multimodal input using Deformable CNN. As the goal of deformable convolutions is to use spatial adaptation of kernels, we only use the operation once, in the last layer. In this way, the most relevant features extracted from RGB are simply enhanced by offsets based on depth information. To this end, two techniques were utilized for calculating the convolutional offsets for the deformable layer.
@@
(iii-b) The second technique, which we refer to as ZACN, is based on [wu2020depthadaptedcnnrgbdcameras]. The intrinsic camera parameters used by this method were obtained from the data collection device.